### v1

# The integral form of Gauss’ law for a generic case of the electrodynamics

Abstract
We have analytically examined the integral form of Gauss’ law for arbitrarily moving charges inside and outside an arbitrarily moving and deforming Gaussian surface with the Maxwell equations under consideration. We have obtained a simple looking differential equation $\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}$ for time-dependent Gauss’ flux-integral. We have explicitly calculated the time-derivative of the flux-integral for the electric field passing through the Gaussian surface for such a generic non-static case and have found no changes in the form of Gauss’ law.
pacs:
01.55.+b General physics, 03.50.De Classical electromagnetism, Maxwell equations, 01.40.Fk Research in physics education
Classical electrodynamics is a well-established branch of physics and engineering since its one and half centuries old (1865 AD) formulation with the Maxwell equations Born . While the integral form of Gauss’ law (1835 AD) is a pillar for electrostatics, the differential form of Gauss’ law is a pillar for the electrodynamics because it is one of the four Maxwell equations Born . This article, however, is related to the formulation of the classical electrodynamics because we are going to examine the validity of the integral form of Gauss’ law for a generic time-dependent case with the Maxwell equations under consideration.
Let us follow the usual notation in the field such as – $\vec{E}=\vec{E}(\vec{r},t)$ ($\vec{B}=\vec{B}(\vec{r},t)$) is the electric field (magnetic field) at the position $\vec{r}$ at time $t$ in the 3-D Euclidean (free) space where $\rho(\vec{r},t)$ and $\vec{J}(\vec{r},t)$ represent the charge density and the current density, respectively. The Maxwell equations for the electromagnetic field in electrodynamics are as follows:
$\divergence\vec{E}(\vec{r},t)=\frac{\rho(\vec{r},t)}{\epsilon_{0}},$ (1)
×$\curl\vec{E}(\vec{r},t)=-\frac{\partial\vec{B}(\vec{r},t)}{\partial t},$ (2)
$\divergence\vec{B}(\vec{r},t)=0,~{}\text{and}$ (3)
×$\curl\vec{B}(\vec{r},t)=\mu_{0}\vec{J}(\vec{r},t)+\mu_{0}\epsilon_{0}\frac{\partial\vec{E}(\vec{r},t)}{\partial t}.$ (4)
Here-from one can easily show that the electromagnetic wave (light) propagates with the speed $c=1/\sqrt{\mu_{0}\epsilon_{0}}$ in the free space Born .
Figure 1: A schematic diagram for $i=1,2,3,4,5$, and $6$ with $n=6$. The (point) charges $q_{2}$, $q_{3}$, $q_{4}$, and $q_{5}$ are moving with velocities $\vec{v}_{2}$, $\vec{v}_{3}$, $\vec{v}_{4}$, and $\vec{v}_{5}$, respectively, inside the moving and deforming Gaussian surface $s(t)$ at time $t$. The surface becomes $s(t+\delta t)$ at time $t+\delta t$. While the charge $q_{1}$ is about to leave $s(t)$ with velocity $\vec{v}_{1}$, the charge $q_{6}$ which although is at rest, is about to come inside $s(t)$ due to motion of the surface, say, with the velocity $\vec{v}^{\prime}$ at the point $\vec{r}^{\prime}$ at time $t$.
While the differential form of Gauss’ law is given by Eqn. (1) even in a non-static case (i.e. $\frac{\partial\rho}{\partial t}\neq 0$ and or $\frac{\partial\vec{J}}{\partial t}\neq 0$), the integral form of Gauss’ law (Gauss’ flux theorem) is given by the following form
$\oint_{s}\vec{E}(\vec{r}^{\prime})\cdot\text{d}\vec{s}=\frac{q_{\text{in}}}{\epsilon_{0}}$ (5)
in the standard textbooks Jackson ; Griffiths for a static case (i.e. for $\frac{\partial\rho}{\partial t}=0$ and $\frac{\partial\vec{J}}{\partial t}=0$) of the total change $q_{\text{in}}$ inside a (closed) Gaussian surface $s$ where a point is denoted by $\vec{r}^{\prime}$. Laws of electrostatics and magnetostatics such as Biot-Savart law and Coulomb’s law, respectively, were already generalized for the electrodynamics with the time-dependence under consideration Griffiths2 . The differential form of Gauss’ law, i.e. Eqn. (1), on the other hand, is valid for a generic case of electrodynamics Born . Thus, a natural question comes - whether the differential form of Gauss’ law, i.e. Eqn. (5), is also valid for a generic non-static (dynamic) case of the charges moving arbitrarily inside and outside an arbitrarily moving and deforming Gaussian surface. A special case of such a study with a fixed amount of charges inside the moving and deforming surface, however, was discussed before Rothwell . Two even more special cases of such a study were also separately discussed (i) for a single charge uniformly moving and (ii) for an accelerated radiating charge, inside a static Gaussian surface under the consideration of the retarded scalar potential Garcia .
The left-hand side of Eqn. (5) represents the flux of the electric field passing through the Gaussian surface $s$. For a non-static case as depicted in figure 1, the flux of the electric field passing through a moving Gaussian surface $s(t)$ can be given by the time-dependent integral $\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)$. This flux-integral may vary with time if the Gaussian surface element $\text{d}\vec{s}(t)$ at position $\vec{r}^{\prime}$ and time $t$ moves with a velocity $\vec{v}^{\prime}(\vec{r}^{\prime},t)$ and the (point) charges $\{q_{i}\}$ ($i=0,1,2,3,...,n$) both inside and outside $s(t)$, as depicted in figure 1, move with the velocities $\{\vec{v}_{i}(\vec{r}_{i},t)\}$. By notation, we have $\vec{r}=\vec{r}^{\prime}$ for $\vec{r}$ on the surface $s(t)$. Otherwise, $\vec{r}$ denotes a point inside and outside $s(t)$. The time-derivative of the flux-integral can be obtained by following the procedure described in Ref.Jackson1 ; Zangwill as
$\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)$ $=$ $\oint_{s(t)}\text{d}\vec{s}(t)\cdot\bigg{[}\frac{\partial\vec{E}(\vec{r},t)}{\partial t}+\vec{v}^{\prime}(\vec{r},t)($
$\divergence\vec{E})-\curl{[\vec{v}^{\prime}(\vec{r},t)\times\vec{E}(\vec{r},t)]}\bigg{]}.$
Now, using the Maxwell equations, especially Eqns. (1) and (4), we recast Eqn. (The integral form of Gauss’ law for a generic case of the electrodynamics) as
$\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)$ $=$ $c^{2}\oint_{s(t)}\text{d}\vec{s}(t)\cdot\curl\bigg[\vec{B}-\frac{\vec{v}^{\prime}\times\vec{E}}{c^{2}}\bigg{missing}]$ (7)
$+\frac{1}{\epsilon_{0}}\oint_{s(t)}\text{d}\vec{s}(t)\cdot\big{[}\rho\vec{v}^{\prime}-\vec{J}\big{]}.$
It is to be noted that $\vec{J}\neq\rho\vec{v}^{\prime}$ because $\vec{v}^{\prime}$ does not represent the velocity of a charge, rather it represents the velocity of the Gaussian surface. Let us now use Gauss’ divergence theorem to the 1st term on the right-hand side of Eqn. (7) to get the following
$\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)$ $=$ $c^{2}\int_{V(t)}\text{d}^{3}\vec{r}\divergence{\bigg{\{}\curl\bigg[\vec{B}-\frac{\vec{v}^{\prime}\times\vec{E}}{c^{2}}\bigg{missing}]\bigg{\}}}$ (8)
$+\frac{1}{\epsilon_{0}}\oint_{s(t)}\text{d}\vec{s}(t)\cdot\big{[}\rho\vec{v}^{\prime}-\vec{J}\big{]}$
where $V(t)$ is the volume of the free space inside the closed surface $s(t)$. Interestingly, the 1st term on the right-hand side of Eqn. (8) is zero as the divergence of the curl of a vector field is identically zero. Now, it is to be noted that $\vec{J}(\vec{r}^{\prime},t)-\rho\vec{v}^{\prime}(\vec{r}^{\prime},t)$ is the net outward current density with respect to the moving surface at the point $\vec{r}^{\prime}$ at time $t$. Thus, the integral on the 2nd term on the right-hand side of Eqn. (8) is nothing but the net current ($I^{(s)}_{\text{in}}(t)=\oint_{s(t)}\text{d}\vec{s}(t)\cdot\big{[}\rho\vec{v}^{\prime}-\vec{J}]$) flowing into the entire moving surface at time $t$. Thus, we recast Eqn. (8) as
$\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}.$ (9)
This is the equation for time-dependent Gauss’ flux-integral. This equation is simple-looking, novel, and useful to determine the flux-integral at any instant of time. Now, integrating both the sides of Eqn. (9) with respect to time, we get
$\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)-\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})=\frac{\triangle q_{\text{in}}(t,t_{0})}{\epsilon_{0}}$ (10)
where $\triangle q_{\text{in}}^{(s)}(t,t_{0})=\int_{t_{0}}^{t}I^{(s)}_{\text{in}}(t^{\prime})\text{d}t^{\prime}$ is the net charge accumulated inside the Gaussian surface $s(t)$ through the entire moving surface $s(t^{\prime})$ in the course of time for $t_{0}<t^{\prime}<t$. Let us now suitably choose the reference time $t_{0}$ such that both the charges and the Gaussian surface were static from $t=-\infty$ to $t=t_{0}$ with the total amount of charge $q_{in}^{(0)}$ in it. Thus, the flux-integral $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})$ follows Gauss’ law $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})=\frac{q_{\text{in}}^{(0)}}{\epsilon_{0}}$ as mentioned in Eqn. (5) for the electrostatics Jackson ; Griffiths . Thus, we recast Eqn. (10) as
$\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{q_{\text{in}}(t)}{\epsilon_{0}}$ (11)
where $q_{\text{in}}(t)=q_{\text{in}}^{(0)}+\triangle q_{\text{in}}^{(s)}(t,t_{0})$ is the net charge inside the Gaussian surface $s(t)$ at any instant of time $t$ for $t>t_{0}$. Eqn. (11) is the desired result for a generic non-static case of charges moving inside and outside an arbitrarily moving and deforming Gaussian surface. It is clear from Eqn. (11) and Eqn. (5) that the integral form of Gauss’ law remains unaltered, i.e. robust, even for a non-static case with a moving and deforming Gaussian surface. The total amount of charges inside the moving surface $s(t^{\prime})$ (for $t_{0}<t^{\prime}<t$) was not necessarily considered to be conserved for the derivation of Eqn. (11). The total charge inside the moving surface would have been conserved if we had considered $I^{(s)}_{\text{in}}(t)=0$ for $t_{0}<t^{\prime}<t$.
Let us now consider a special case of the conservation of the total charge inside the moving Gaussian surface. It is interesting to note from Eqn. (11) that, if the charges arbitrarily move inside the arbitrarily moving and deforming surface $s(t^{\prime})$ and no current passes through it for $t_{0}<t^{\prime}<t$, then we have $\triangle q_{\text{in}}^{(s)}(t,t_{0})=0$ for $t_{0}<t^{\prime}<t$. In such a special case, Eqn. (11) takes the form
$\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{q^{(0)}_{\text{in}}}{\epsilon_{0}}.$ (12)
It should be mentioned that the flux of the electric field through a part of the moving Gaussian surface $s(t^{\prime})$ though changes in time for $t_{0}<t^{\prime}<t$, the flux through whole the Gaussian surface, of course, does not change at all in such a special case. This special case [^1] The case was made special in Ref.Rothwell for the use of the continuity equation for the charges inside the moving surface. was studied in Ref.Rothwell without an explicit derivation. Eqn. (12) is the same as that obtained in Ref.Rothwell . Eqn. (12) is also compatible with the previous study for the even more special cases of the uniform (relativistic) motion of a charge inside a static Gaussian surface and the accelerated motion of a radiating charge inside a static Gaussian surface Garcia .
To conclude, we have obtained a simple looking differential equation for time-dependent Gauss’ flux-integral for charges arbitrarily moving inside and outside an arbitrarily moving and deforming Gaussian surface. There-from we have explicitly calculated the time-dependent flux-integral for the electric field passing through the Gaussian surface and have found robustness in the original form of Gauss’ law.
Our calculations are based on the Maxwell equations which don’t need any relativistic correction. Hence our final result (i.e. Eqn. (11)) on the time-dependent form of Gauss’ law is essentially correct from the relativistic point of view.
The charges inside and outside the arbitrarily moving and deforming Gaussian surface are considered to be moving arbitrarily. While the uniformly moving charges result in no radiation, the accelerating charges inside and outside the Gaussian surface result in electromagnetic radiation. This radiation does not change the form of Gauss’ law as clear from Eqn. (11).
Finally, it should be mentioned that the differential equation for Gauss’ flux-integral (Eqn. (9)) obtained by us, is novel, simple looking, and important for determining time-dependent Gauss’ flux-integral. The time-dependent form of Gauss’ flux-integral (Gauss’ law), i.e. Eqn. (11), obtained by us, is also absolutely new. Our explicit derivation of Gauss’ flux-integral for a generic non-static case, however, can be directly useful even in an undergraduate or postgraduate physics class on electrodynamics for a better understanding of Gauss’ law. Gauss’ flux-integral obtained by us would be applicable for arbitrary charge and current distributions with radiating charges and non-conserved total charge in a region.
Data availability statement
No new data were created or analysed in this study.
Acknowledgement
S. Biswas acknowledges the partial financial support of the SERB (now ANRF), DST, Govt. of India under the EMEQ Scheme [No. EEQ/2023/000788]. We thank Mr. Bikram Keshari Behera (UoH, India) for helping us drawing figure 1.